0.5 t-0.05 w, where w is the hogs 2 nts weight (in lbs) by the time its t days old. You just care about the weight (w) of a 30-day old hog 3.3 The growth model of a production hog is given by 3 log,ow Pts (i.e. just at t 30), but the non-linear equation for w is too hard to solve for w directly. Write this problem in the generic f(x) 0 form for the case t 30. Thats it! m just asking you to complete the physics step-dont actually solve for w! Just write the initial f(x) equation, ready to solve for the hogs weight using any of the Chapter 3 methods later.
Answer #1

3*log10w = 0.5*t - 0.05*w

putting , t= 30

3*log10w = 0.5*30 - 0.05*w

3*log10w = 15 - 0.05*w

log10w3 = 15- 0.05*w

log10w3 + 0.05*w -15 =0